# Real-time Bayesian inference at extreme scale:

### v1

@@ VI How Performance Was Measured
Our inference and prediction framework consists of two code bases. The first is the “Cascadia application code,” a C++ solver for the acoustic–gravity model (1) implemented in MFEM. This code is used in the (offline) Phase 1 of our framework to compute adjoint PDE solutions that define the p2o map $\mathbf{F}$ and p2q map $\mathbf{F}_{\hskip-1.0ptq}$. Both AMD (via HIP) and NVIDIA (via CUDA) GPU-based systems are supported; we report performance results for both types of systems.
The second code, which implements Phases 2–4 of our framework, is an FFT-based C++/CUDA code that exploits the block Toeplitz structure of $\mathbf{F}$ and $\mathbf{F}_{\hskip-1.0ptq}$, together with the other algorithmic innovations described in the preceding section, to solve the real-time Bayesian parameter inference and QoI prediction problems. This code makes extensive use of CUDA-exclusive libraries, including cuDSS and cuSOLVERMp, among others, so we report performance results on NVIDIA GPU-based systems.
@@ VI-A HPC systems
LLNL’s El Capitan is an HPE Cray EX supercomputer with 11,136 nodes, each containing 4 AMD Instinct™ MI300A Accelerated Processing Units (APUs). The MI300A consists of 24 Zen4-based CPU cores, a CDNA3-based GPU, and 128 GB of HBM3 memory, integrated onto a single package. Each MI300A has a peak double precision throughput of 61.3 TFLOP/s for a total peak of the machine of 2.73 EFLOP/s. The system nodes are connected by an HPE Slingshot-200 dragonfly interconnect with at most three network hops between any two nodes.
NERSC’s Perlmutter is an HPE Cray Shasta supercomputer with 1,536 nodes, each containing an AMD EPYC™ 7763 64-core CPU and 4 NVIDIA A100 GPUs. Each GPU has 40 GB of HBM2 memory and a peak double precision throughput of 9.7 TFLOP/s for a total system peak of 59.6 PFLOP/s. The system is connected by a Slingshot-11 dragonfly interconnect with at most three network hops between any two nodes.
@@ VI-B MFEM library
MFEM is a C++ finite element (FE) library that provides high-performance discretization algorithms to a wide range of HPC application codes in national laboratories, industry, and academia [anderson2021mfem]. In addition to SoA support for high-order methods, MFEM has been designed to be highly scalable on the full range of computing platforms, from laptops to GPU-accelerated supercomputers [mfem2024]. Building on efforts in the U.S. Exascale Computing Project [ceed], the library has put a special emphasis on GPU architectures. An important algorithmic innovation that enables MFEM to achieve high performance on GPU hardware is the concept of FE operator decomposition that exposes data parallelism in high-order FE computations [ceed2020scalability], and enables additional optimizations on quadrilateral and hexahedral grids by taking advantage of the Cartesian product structure at the element level in the decomposition. Critically, MFEM’s approach also leads to a significantly reduced memory requirement through partial assembly (PA), which stores an asymptotically optimal amount of data: $\mathcal{O}$(1) per degree of freedom (DOF). Compared to the classical full (global sparse matrix) or element-level (local dense matrices) assembly algorithms, PA leads to orders-of-magnitude faster execution times on both CPU and GPU architectures. Another option supported by MFEM is matrix-free (MF) assembly where no data is stored and all computations are done on the fly; see Fig. 7. In this work, we focus on faster time-to-solution (as opposed to higher FLOP/s), so we use the PA approach to discretize the forward problem (1).
@@ VI-C Cascadia application code
where $\vec{u},\vec{\tau}\in(L^{2}(\Omega))^{3}$ and $p,v\in H^{1}(\Omega)$; (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes the (component-wise) $L^{2}(\Omega)$ inner product, and $\langle\cdot,\cdot\rangle_{\partial\Omega}$ is the $L^{2}(\partial\Omega)$ inner product over (part of) the boundary $\partial\Omega$.
TABLE II: Scalability setup on the El Capitan and Perlmutter systems
@@
Nodes GPUs Processor grid Elements Elements/GPU
@@
170 680 10 $\times$   17 $\times$ 4 3,386,900,480 2,490,368
340 1,360 10 $\times$   34 $\times$ 4 6,773,800,960 1,245,184
680 2,720 20 $\times$   34 $\times$ 4 13,547,601,920 622,592
1,360 5,440 20 $\times$   68 $\times$ 4 27,095,203,840 311,296
2,720 10,880 40 $\times$   68 $\times$ 4 54,190,407,680 155,648
5,440 21,760 40 $\times$ 136 $\times$ 4 108,380,815,360 77,824
@@
94 376 2 $\times$   47 $\times$ 4 591,396,864 786,432
188 752 2 $\times$   94 $\times$ 4 1,182,793,728 393,216
376 1,504 4 $\times$   94 $\times$ 4 2,365,587,456 196,608
752 3,008 4 $\times$ 188 $\times$ 4 4,731,174,912 98,304
@@
The primary performance metric that matters from our application point of view is therefore runtime per timestep, which we report in our weak and strong scalability studies. Reported runtimes were averaged over ten consecutive timesteps (40 operator applications), discarding the initial ten timesteps as warm-up. The problem sizes we used to test scalability on the El Capitan and Perlmutter systems are summarized in Table II. Since runtime per timestep is the primary application-relevant performance metric, we optimized GPU-kernel performance for DOF throughput rather than total achieved FLOP/s (see §VII-B); FLOP counts and memory access rates on El Capitan were analyzed using the ROCm Compute profiler.
@@ VI-D FFT-based Bayesian inference code
The FFT-based algorithm used to enable fast Hessian matvecs is completely agnostic to the particular application; it is a general purpose code that can be used to calculate matvecs with any block-triangular Toeplitz matrix. Examples of other digital twin applications that fit into this framework are discussed in §VIII. This code employs the cuFFT and cuBLAS libraries to compute batched FFTs and batched dense matvecs, respectively. Performance of the matvecs was measured with the NSight Compute profiler. The primary performance goal of this code is to achieve fast time-to-solution. FFT matvec wall timings were measured with omp_get_wtime after calling cudaDeviceSynchronize and MPI_Barrier. The implementation of the FFT matvec algorithm is available open-source at https://github.com/s769/FFTMatvec.
The prior solves used to form the $\mathbf{G}^{*}$ matrix were implemented with the cuDSS library. Linear solves with the dense, symmetric $\mathbf{K}$ matrix were implemented via Cholesky factorization through the cuSOLVERMp library.
@@ VII-A Scalability
Figure 5: Weak scalability (left) and strong scalability (right) results on El Capitan, from 85 nodes (340 AMD MI300A GPUs) to 10,880 nodes (43,520 GPUs), and on Perlmutter, from 47 nodes (188 NVIDIA A100 GPUs) to 1,504 nodes (6,016 GPUs). Numbers along the graph lines indicate parallel efficiency.
We present weak and strong scalability results (Fig. 5) on the El Capitan and Perlmutter supercomputers. The computational cost of our inference and prediction framework is overwhelmingly dominated by the cost of the adjoint wave propagation PDE solutions in Phase 1 (offline) that are used to construct the p2o map $\mathbf{F}$ and p2q map $\mathbf{F}_{\hskip-1.0ptq}$. The computational cost of each adjoint PDE solution, in turn, is overwhelmingly dominated by the RK4 time-stepping solver; each time step involves four applications of the PDE operator. For this offline phase, scalability to large problem sizes is critical (in addition to throughput), since we wish to apply our framework to other settings beyond Cascadia where tsunamis can impact populations across long distances (such as occurred in the 2004 Sumatra–Andaman earthquake).
On El Capitan, the solver maintained 92% weak parallel efficiency over a 128-fold increase in problem size, from 85 nodes (340 AMD MI300A GPUs) to 10,880 nodes (43,520 GPUs) of the full system. The largest problem involved over 55.5 trillion DOF, which, to the best of our knowledge, is the largest reported unstructured FE computation. This computation used about 86% (4.8 PB) of the total available machine memory (5.57 PB). For the largest problem fitting on 340 GPUs (434 billion DOF), the solver achieved a parallel speedup of 100.9 over a 128-fold increase in GPUs, a strong parallel efficiency of 79%.
We also obtained excellent weak and strong scalability on the Perlmutter system over a 32-fold increase in GPUs, from 47 nodes (188 NVIDIA A100 GPUs) to 1,504 nodes (6,016 GPUs). Here, the solver achieved 100% weak efficiency for a problem with 403 million DOF per GPU and 92% strong efficiency for a problem involving 75.7 billion DOF.
While GPU-based systems have been the focus of our performance optimizations, we want to emphasize that our solver also exhibits excellent weak and strong scalability on CPU-based systems. For example, on TACC’s Frontera [^2] Frontera hosts 8,368 Intel Xeon Platinum 8280 (“Cascade Lake”) compute nodes contained in 101 racks. Each node has 56 cores on two sockets and 192 GB of DDR4 RAM; see https://docs.tacc.utexas.edu/hpc/frontera/#system. supercomputer, we achieve an outstanding weak efficiency of 95% over a 8,192-fold increase in problem size, from 1 node (56 cores) to 8,192 nodes (458,752 cores), with the largest problem involving 2.20 trillion DOF (4.80 million DOF per core), and a strong efficiency of 70% over a 128-fold increase in cores, from 3,584 cores to 458,752 cores.
@@ VII-B Peak performance
Figure 7: Throughput, in billion degrees of freedom (GDOF) per second, for the kernels corresponding to the off-diagonal blocks in (4) on a single GPU of El Capitan. Measured FLOP/s rates with percentage of peak (61.3 TFLOP/s) are shown in the saturated regime for each curve. By utilizing the GPU shared memory we improved the initial partial assembly (PA) kernels (red) to a version (brown) that is 13$\times$ faster. Specifying explicit launch bounds for the kernels allowed us to additionally improve performance (yellow), which was the version of the kernels used in the scaling runs in Fig. 5. Further optimizations by fusing the actions of the two operators into one kernel allowed us to achieve a peak performance of 24 GDOF/s (blue) corresponding to 4 TFLOP/s, or 6.51% of FLOP/s peak. Although the matrix-free (MF) approach (green) attains higher FLOP/s, its time-to-solution (DOF throughput) remains lower than that of Fused PA.
We also performed extensive optimization of the memory usage, including: instrumenting the code to track memory usage separately on host and device; taking advantage of the unified memory on the MI300A APU by freeing host allocation so they can be used for GPU allocations; using sparsity in the RHS to avoid storing full vectors; fusing permutation operators with operator action kernels; recomputing determinants of Jacobian matrices instead of storing them; batching multiple small memory allocations together; switching to 64-bit integers for certain variables local to the GPU (e.g., multi-dimensional quadrature point data); and carefully reusing temporary vectors from RK4 for temporary vectors needed in the operator action. With all these optimizations for a problem of size 67 million DOF per APU, we improved unified memory usage from (5.2 host + 30.7 device) to (1.1 host + 5.64 device) GiB/APU, a reduction of 5.33$\times$. These improvements allowed us to fit $\sim$1.28 billion DOF per APU.
Our primary performance metric is time-to-solution, which, for the PDE operator (4), is measured as DOF throughput. For a given problem size, higher throughput is directly proportional to faster computation time. As shown in Fig. 7, more FLOP/s do not imply faster time-to-solution: the best-performing implementation, Fused PA, has a 1.9×1.9$\times$1.9 × lower percentage of FLOP/s peak compared to Fused MF, but is 1.6×1.6$\times$1.6 × faster. The reason for this is that although Fused MF requires less memory, it incurs a higher computational cost, with an arithmetic intensity of 17.2 FLOP/byte compared to Fused PA’s 2.5 FLOP/byte. Although the percentage of peak FLOP/s in Fig. 7 is not as high as for dense matrix operations, we note that it is much better than the results reported for the HPCG benchmark on comparable systems [^3] As of April 15, 2025, HPCG results have not been released for El Capitan. (e.g., 0.7% of peak of Frontier’s MI250Xs). Similar to the type of computations performed in FE algorithms, HPCG has a FLOP count and memory access rate proportional to the number of DOF.
All GPU kernels involved in the FFT matvec implementation (§VI-D) are memory-bound and achieve 80–95% of peak memory bandwidth on NVIDIA A100 and GH200 GPUs [venkat2024fft].
VII-C Time-to-solution
We measured time-to-solution for a margin-wide rupture scenario in the CSZ for which we presented inversion results in §V. These computations were performed on 128 Perlmutter nodes (512 NVIDIA A100 GPUs) and involved inference of more than one billion parameters from 252,000 synthetic, noisy observations obtained by 600 hypothesized seafloor pressure sensors, followed by real-time forecasting of the QoI (tsunami wave heights) under uncertainty for 21 QoI locations.
For this problem, Table III summarizes the compute times for each phase of our inference and prediction framework. As discussed above, the total computational expense is dominated by the PDE solutions in the (offline) Phase 1; in these offline computations, scalability is the most important metric, whereas time-to-solution is the metric that matters most for the Phase 4 (online) computations that are executed in real time.
Phase Task GPUs Compute time
@@ VII-C Time-to-solution
factorize $\mathbf{K}$ 25 22 s
[…]
4 infer parameters $\mathbf{m}_{\text{map}}$ 512 $<$ 0.2 s
predict QoI $\mathbf{q}_{\text{map}}$ 1 $<$ 1 ms
Phase 1 relies on the PDE solver to precompute the p2o map $\mathbf{F}$ and p2q map $\mathbf{F}_{\hskip-1.0ptq}$. For the given configuration (600 sensors, 21 QoI locations), we performed a total 621 PDE solutions, each of which (on average) took 52 minutes on 512 A100 GPUs, for a total of 538 hours of compute time (note that each one of the 621 PDE solutions can be computed independently).
Once Phase 1 was completed, we exploited the block Toeplitz structure of the precomputed $\mathbf{F}$ and $\mathbf{F}_{\hskip-1.0ptq}$ to perform fast FFT-based Hessian matvecs. Each Hessian matvec, conventionally requiring a pair of forward and adjoint PDE solutions that take 104 minutes on 512 A100 GPUs per matvec, could now be performed in 0.024 seconds (averaged over 100 matvecs) on the same number of GPUs, a speedup of 260,000. We re-emphasize that these FFT-based Hessian matvecs are exact and do not incorporate any approximations. The 260,000$\times$ speedup derives primarily from the parallel algorithmic innovations, discussed in §V, that massively reduce the overall complexity of the Hessian matvec which, after the precomputations in Phase 1, avoids PDE solutions altogether.
These rapid Hessian matvecs are the workhorse for the remaining offline computations (Phases 2–3). For example, forming the data space Hessian $\mathbf{K}$ requires 252,000 matvecs—an intractable task if each matvec required a forward/adjoint pair of PDE solutions—which we computed with the FFT-based matvecs in just 100 minutes. Finally, real-time inference of the parameters (seafloor motion) and forward prediction of the QoI (surface wave heights) was achieved in less than 0.2 seconds of compute time on 512 A100 GPUs.
We conclude by noting that computing the inverse solution with SoA methods (see §IV), without the innovations described in §V, would not only have been prohibitive in real time, but would have been intractable (50 years on 512 A100 GPUs). With the advances described in this paper, the computing time reduces to a one-time offline computation of 621 ($N_{d}+N_{q}$) adjoint wave propagations, requiring 538 hours on 512 GPUs and representing an $\sim$810$\times$ reduction in PDE solutions. Then, for each earthquake that excites the sensors, the online solution of the Bayesian inverse problem requires $<0.2$ s, a factor of $10^{10}$ speedup over the SoA CG algorithm.

### v2

@@ VI How Performance Was Measured
Our inference and prediction framework consists of two code bases. The first is the “Cascadia application code,” a C++ solver for the acoustic–gravity model (1) implemented in MFEM. This code is used in the (offline) Phase 1 of our framework to compute adjoint PDE solutions that define the p2o map $\mathbf{F}$ and p2q map $\mathbf{F}_{\thinspace q}$. Both AMD (via HIP) and NVIDIA (via CUDA) GPU-based systems are supported; we report performance results for both types of systems.
The second code, implementing Phases 2–4, consists of two parts. The first is “FFTMatvec,” which computes matvecs with the block Toeplitz matrices $\mathbf{F}$ and $\mathbf{F}_{\thinspace q}$ using a portable C++/CUDA/HIP implementation based on §V-A. Performance is reported on AMD and NVIDIA systems. The second part solves the real-time Bayesian parameter inference and QoI prediction problems using the algorithmic innovations in §V. It remains CUDA-exclusive, using libraries like cuDSS and cuSOLVERMp; performance is reported on NVIDIA systems.
@@ VI-A HPC systems
LLNL’s El Capitan is an HPE Cray EX supercomputer with 11,136 nodes, each containing 4 AMD Instinct MI300A Accelerated Processing Units (APUs). The MI300A consists of 24 Zen4-based CPU cores, a CDNA3-based GPU, and 128 GB of HBM3 memory, integrated onto a single package. Each MI300A has a 61.3 TFLOP/s peak double precision throughput for a total machine peak of 2.73 EFLOP/s. The system is connected by an HPE Slingshot-200 dragonfly interconnect with at most three network hops between any two nodes.
CSCS’s Alps is an HPE Cray EX supercomputer with 2,688 nodes, each containing 4 NVIDIA GH200 superchips. Each GH200 integrates a Grace CPU (72 Arm cores) and an H100 GPU (96 GB HBM3, 34.0 TFLOP/s peak double precision throughput); the system peak is 574.8 PFLOP/s. The system is connected by an HPE Slingshot-11 dragonfly interconnect.
NERSC’s Perlmutter is an HPE Cray Shasta supercomputer with 1,536 nodes, each containing an AMD EPYC 7763 64-core CPU and 4 NVIDIA A100 GPUs. Each GPU has 40 GB of HBM2 memory and a peak double precision throughput of 9.7 TFLOP/s for a total system peak of 59.6 PFLOP/s. The system is connected by a Slingshot-11 dragonfly interconnect.
@@ VI-B MFEM library
MFEM is a C++ finite element (FE) library that provides high-performance discretization algorithms to a wide range of HPC application codes in national laboratories, industry, and academia [29]. In addition to SoA support for high-order methods, MFEM has been designed to be highly scalable on the full range of computing platforms, from laptops to GPU-accelerated supercomputers [30]. Building on efforts in the U.S. Exascale Computing Project [31], the library has put a special emphasis on GPU architectures. An important algorithmic innovation that enables MFEM to achieve high performance on GPU hardware is the concept of FE operator decomposition that exposes data parallelism in high-order FE computations [32], and enables additional optimizations on quadrilateral and hexahedral grids by taking advantage of the Cartesian product structure at the element level in the decomposition. Critically, MFEM’s approach also leads to a significantly reduced memory requirement through partial assembly (PA), which stores an asymptotically optimal amount of data: $\mathcal{O}$(1) per degree of freedom (DOF). Compared to the classical full (global sparse matrix) or element-level (local dense matrices) assembly algorithms, PA leads to orders-of-magnitude faster execution times on both CPU and GPU architectures. Another option supported by MFEM is matrix-free (MF) assembly where no data is stored and all computations are done on the fly; see Fig. 7. In this work, we focus on faster time-to-solution (as opposed to higher FLOP/s), so we use the PA approach to discretize the forward problem (1).
@@ VI-C Cascadia application code
where $\vec{u},\vec{\tau}\in(L^{2}(\Omega))^{3}$ and $p,v\in H^{1}(\Omega)$; $(\cdot,\cdot)$ denotes the (component-wise) $L^{2}(\Omega)$ inner product, and $\langle\cdot,\cdot\rangle_{\partial\Omega}$ is the $L^{2}(\partial\Omega)$ inner product over (part of) the boundary $\partial\Omega$.
TABLE II: Scalability setup on El Capitan, Alps and Perlmutter
@@
Nodes GPUs Processor grid Mesh elements Elements/GPU
@@
The primary performance metric that matters from our application point of view is therefore runtime per timestep, which we report, together with parallel efficiency and relative speedup, in our weak and strong scalability studies. Reported runtimes were averaged over ten consecutive timesteps (40 operator applications), discarding the initial ten timesteps as warm-up. The mesh sizes we used to test scalability on El Capitan, Alps, and Perlmutter are summarized in Table II. Since runtime per timestep is the primary application-relevant performance metric, we optimized GPU-kernel performance for DOF throughput rather than total achieved FLOP/s (see §VII-B); FLOP and byte counts were manually calculated, and memory access rates were analyzed using the ROCm profiler on El Capitan and Nsight Compute on Alps.
@@ VI-D FFT-based Bayesian inference code
The performance-portable, open-source FFTMatvec [^2] https://github.com/s769/FFTMatvec is application agnostic and can be used with any block-triangular Toeplitz matrix; see §VIII. It uses vendor libraries (e.g., {cu,roc}FFT, {cu,roc}BLAS) for batched computations. Its performance was profiled with Nsight Compute and rocprofv3.
The CUDA-based Bayesian inference code uses cuDSS for prior solves to form the $\mathbf{G}^{*}$ matrix and cuSOLVERMp for linear solves with the dense, symmetric $\mathbf{K}$ matrix via Cholesky factorization. Wall timings for FFTMatvec and the Bayesian inference codes were measured with omp_get_wtime after device synchronization and MPI_Barrier. The primary performance goal of both codes is a fast time-to-solution.
@@ VII-A Scalability
Figure 5: Weak scalability (left) and strong scalability (right) results on El Capitan, from 85 nodes (340 AMD MI300A GPUs) to 10,880 nodes (43,520 GPUs), on Alps, from 36 nodes (144 NVIDIA GH200 GPUs) to 2,304 nodes (9,216 GPUs), and on Perlmutter, from 47 nodes (188 NVIDIA A100 GPUs) to 1,504 nodes (6,016 GPUs). Numbers along the graph lines indicate parallel efficiency.
We present weak and strong scalability results (Fig. 5) on the El Capitan, Alps, and Perlmutter supercomputers. The computational cost of our inference and prediction framework is overwhelmingly dominated by the cost of the adjoint wave propagation PDE solutions in Phase 1 (offline) that are used to construct the p2o map $\mathbf{F}$ and p2q map $\mathbf{F}_{\thinspace q}$. The computational cost of each adjoint PDE solution, in turn, is overwhelmingly dominated by the RK4 time-stepping solver; each timestep involves four applications of the PDE operator. Scalability to large problem sizes is critical (in addition to throughput) for this offline phase, since we wish to apply our framework to other settings beyond Cascadia where tsunamis can impact populations across long distances (such as occurred in the 2004 Sumatra–Andaman earthquake).
On El Capitan, the solver maintained 92% weak parallel efficiency over a 128-fold increase in problem size, from 85 nodes (340 AMD MI300A GPUs) to 10,880 nodes (43,520 GPUs). The largest problem involved over 55.5 trillion DOF, which, to the best of our knowledge, is the largest reported unstructured mesh FE computation. This computation used 86% (4.8 PB) of the available machine memory (5.57 PB). For the largest problem fitting on 340 GPUs (434 billion DOF), the solver achieved a parallel speedup of 100.9 over a 128-fold increase in GPUs—a strong parallel efficiency of 79%.
We also obtained excellent weak and strong scalability on Alps over a 64-fold increase in GPUs, from 36 nodes (144 NVIDIA GH200 GPUs) to 2,304 nodes (9,216 GPUs). Here, the solver achieved 99% weak parallel efficiency with 1.01 billion DOF per GPU and 91% strong parallel efficiency for a problem involving 145 billion DOF. Similar scalability was observed on Perlmutter over a 32-fold increase in GPUs.
While GPU-based systems have been the focus of our performance optimizations, we want to emphasize that our solver also exhibits excellent scalability on CPU-based systems. On TACC’s Frontera [^3] Frontera hosts 8,368 Intel Xeon Platinum 8280 (“Cascade Lake”) compute nodes contained in 101 racks. Each node has 56 cores on two sockets and 192 GB of DDR4 RAM; see https://docs.tacc.utexas.edu/hpc/frontera/#system. supercomputer, we achieve an outstanding weak parallel efficiency of 95% over an 8,192-fold increase in problem size, from 1 node (56 cores) to 8,192 nodes (458,752 cores), with the largest problem involving 2.20 trillion DOF (4.80 million DOF per core), and a strong parallel efficiency of 70% over a 128-fold increase in cores, from 3,584 cores to 458,752 cores.
@@ VII-B Peak performance
All GPU kernels in FFTMatvec (§VI-D) are memory-bound, achieving 80–95% of peak memory bandwidth on NVIDIA (A100, GH200) and AMD (MI250X, MI300X) GPUs [26].
VII-C Time-to-solution
We measured time-to-solution for a margin-wide rupture scenario in the CSZ for which we presented inversion results in §V. These computations were performed on 128 Perlmutter nodes (512 NVIDIA A100 GPUs) and involved inference of more than one billion parameters from 252,000 synthetic, noisy observations obtained by 600 hypothesized seafloor pressure sensors, followed by real-time forecasting of the QoI (surface wave heights) under uncertainty for 21 QoI locations.
For this problem, Table III summarizes the compute times for each phase of our inference and prediction framework. As discussed above, the total computational expense is dominated by the PDE solutions in the (offline) Phase 1; scalability is the most important metric in these offline computations, whereas time-to-solution is the metric that matters most for the Phase 4 (online) computations that are executed in real time.
Phase Task GPUs Compute time
1 form $\mathbf{F}:\mathbf{m}\mapsto\mathbf{d}$ 512 600 $\times\!$ 52 m $\sim$ 520 h
form $\mathbf{F}_{\thinspace q}:\mathbf{m}\mapsto\mathbf{q}$ 512 21 $\times\!$ 52 m $\sim$ 18 h
2 form $\mathbf{G}^{*}\coloneq\mathbf{\Gamma}_{\thinspace\text{prior}}\mathbf{F}^{*}$ 16 600 $\times\!$ 4.5 s $\sim$ 45 m
form $\mathbf{G}_{q}^{*}\coloneq\mathbf{\Gamma}_{\thinspace\text{prior}}\mathbf{F}_{\thinspace q}^{*}$ 16 21 $\times\!$ 4.5 s $\sim$ 1.5 m
form $\mathbf{K}\coloneq\mathbf{\Gamma}_{\thinspace\text{noise}}+\mathbf{F}\mathbf{G}^{*}$ 512 252k $\times\!$ 24 ms $\sim$ 100 m
factorize $\mathbf{K}$ 25 22 s
3 compute $\mathbf{\Gamma}_{\thinspace\text{post}(q)}$ 512 8,820 $\times\!$ 150 ms $\sim$ 25 m
compute $\mathbf{Q}:\mathbf{d}\mapsto\mathbf{q}$ 512 8,820 $\times\!$ 150 ms $\sim$ 25 m
4 infer parameters $\mathbf{m}_{\text{map}}$ 512 << 0.2 s
predict QoI $\mathbf{q}_{\text{map}}$ 1 << 1 ms
Phase 1 relies on the PDE solver to precompute the p2o map $\mathbf{F}$ and p2q map $\mathbf{F}_{\thinspace q}$. For the given configuration (600 sensors, 21 QoI locations), we performed a total 621 PDE solutions, each averaging 52 minutes on 512 A100 GPUs, for a total of 538 hours of compute time (note that each one of the 621 PDE solutions can be computed independently).
@@ VII-C Time-to-solution
These rapid Hessian matvecs are the workhorse for the remaining offline computations (Phases 2–3). For example, forming the data space Hessian $\mathbf{K}$ requires 252,000 matvecs, an intractable task if each matvec required a forward/adjoint pair of PDE solutions. Instead, it was carried out using our FFT-based matvecs in just 100 minutes. Finally, real-time inference of the parameters (seafloor motion) and forward prediction of the QoI (surface wave heights) was achieved in less than 0.2 seconds of compute time on 512 A100 GPUs.